THE EVERTON COLLEGE STORE PAID $1794 FOR AN ORDER OF 46 CALCULATORS. THE STORE PAID $9 FOR EACH SCIENTIFIC CALCULATOR. THE OTHERS ALL GRAPHING CALCULATORS, COST $55 EACH HOW MANY OF EACH TYPE OF CALCULATOR WAS ORDERED
Answer by mananth(16946)

calculators c numbers , Price = $9.00
graphing calculators g numbers , Price = $55.00
Total 345
1 c + 1 g = 46 .............1
9 c + 55 g = 1794 .............2
Eliminate g
multiply (1)by -55
Multiply (2) by 1
-55 c -55 g = -2530
9 c + 55 g = 1794
Add the two equations
-46 c = -736
/ -46
c = 16
plug value of c in (1)
1 c + 1 g = 46
16 + 1 g = 46
1 g = 46 -16
g = 30

calculators 16 numbers
graphing calculators 30 numbers
m.ananth@hotmail.ca